VIBRATIONAL ENERGY DISPOSAL AND THE SPECTATOR MODEL FOR REACTION OF CH_{3}D WITH Cl

ROBERT J. HOLIDAY, CHAN HO KWON, \underline{CHRISTOPHER J. ANNESLEY}, and F. FLEM-ING CRIM, Department of Chemistry, University of Wisconsin-Madison, Madison, WI 53706.

Product state distributions for the reaction of vibrationally state selected CH_{3}D with Cl, CH_{3}D(ν) + Cl → CH_{2}D + HCl, reveal the extent to which vibrational energy initially deposited in the reactant survives in the CH_{2}D product. Infrared excitation prepares CH_{3}D in either the ν_{1} + ν_{4} (E) or the 2ν_{4} (A_{1}, E) C-H stretching states and resonant enhanced multiphoton ionization (REMPI) detects the CH_{2}D. With ν_{1} + ν_{4} excitation, the reaction forms only vibrationally unexcited products, CH_{2}D (0_{0}), but 2ν_{4} excitation forms CH_{2}D with stretch excited (1_{1}) and stretch-bend combination excited (1_{1}4_{1}) products. These results are consistent with the spectator model. In local mode notation, the ν_{1} + ν_{4} state corresponds to two quanta of excitation in the C-H bond (|200>|0>), and it reacts at the excited bond, giving all ground state CH_{2}D. The 2ν_{4} state corresponds to excitation in two C-H bonds (|110>|0>), and in this case one excited C-H bond breaks while the other excited C-H bond remains in the products, resulting in an C-H stretch excited products.